### v1

@@ Abstract
We propose and analyze a continuous-time robust reinforcement learning framework for optimal stopping problems under ambiguity. In this framework, an agent chooses a stopping rule motivated by two objectives: robust decision-making under ambiguity and learning about the unknown environment. Here, ambiguity refers to considering multiple probability measures dominated by a reference measure, reflecting the agent’s awareness that the reference measure representing her learned belief about the environment would be erroneous. Using the $g$-expectation framework, we reformulate an optimal stopping problem under ambiguity as an entropy-regularized optimal control problem under ambiguity, with Bernoulli distributed controls to incorporate exploration into the stopping rules. We then derive the optimal Bernoulli distributed control characterized by backward stochastic differential equations. Moreover, we establish a policy iteration theorem and implement it as a reinforcement learning algorithm. Numerical experiments demonstrate the convergence and robustness of the proposed algorithm across different levels of ambiguity and exploration.
@@
optimal stopping, ambiguity, robust optimization, $g$-expectation, reinforcement learning, policy iteration.
@@ 1 Introduction
Optimal stopping is a class of decision problems in which one seeks to choose a time to take a certain action so as to maximize an expected reward. It is applied in various fields, for instance to analyze the optimality of the sequential probability ratio test in statistics (e.g., [65]), to study consumption habits in economics (e.g., [18]), and notably to derive American option pricing (e.g., [55]). A common challenge arising in all these fields is finding the best model to describe the underlying process or probability measure, which is usually unknown. Although significant efforts have been made to propose and analyze general stochastic models with improved estimation techniques, a margin of error in estimation inherently exists.
@@
In response to such model misspecification and estimation errors, recent works, Dai et al. [15] and Dong [17], have cast optimal stopping problems within the continuous time reinforcement learning (RL) framework of Wang et al. [66] and Wang and Zhou [67]. Arguably, the exploratory (or randomized) optimal stopping framework is viewed as model-free, since agents, even without knowledge of the true model or underlying dynamics of the environment, can learn from observed data and determine a stopping rule that yields the best outcome. In this sense, the framework provides a systematic way to balance exploration and exploitation in optimal stopping.
However, the model-free view of the exploratory RL framework has a pitfall: the learning environment reflected in observed data often differs from the actual deployment environment (e.g., due to distributional or domain shifts). Consequently, a stopping rule derived from the learning process may fail in practice. Indeed, Chen and Epstein [11] explicitly ask: “Would ambiguity not disappear eventually as the agent learns about her environment?” In response, Epstein and Schneider [22] and Marinacci [42] stress that the link between empirical frequencies (i.e., observed data) and asymptotic beliefs (updated through learning) can be weakened by the degree of ambiguity in the agent’s prior beliefs about the environment. This suggests that ambiguity can persist even with extensive learning, limiting the reliability of a purely model-free framework. Such limitations have been recognized in the RL literature, leading to significant developments in robust RL frameworks such as [9, 45, 48, 59, 69].
@@
The aim of this article is to propose and analyze a continuous-time RL framework for optimal stopping under ambiguity. Our framework starts with revisiting the following optimal stopping problem under $g$-expectation (Coquet et al. [12], Peng [53]): Let ${\mathcal{}T}_{t}$ be the set of all stopping times with values in $[t,T]$. Denote by ${\mathcal{}E}_{t}^{g}[\cdot]$ the (conditional) $g$-expectation with driver $g:\Omega\times[0,T]\times\mathbb{R}^{d}\to\mathbb{R}$ (satisfying certain regularity and integrability conditions; see Definition 2.1), which is a filtration-consistent adverse nonlinear expectation whose representing set of probability measures is dominated by a reference measure $\mathbb{P}$ (see Remark 2.2). Then, the optimal stopping problem under ambiguity is given by
@@
where $(\beta_{t})_{t\in[0,T]}$ is the discount rate, $r:\mathbb{R}^{d}\to\mathbb{R}$ and $R:\mathbb{R}^{d}\to\mathbb{R}$ are reward functions, and $(X_{t}^{x})_{t\in[0,T]}$ is an Itô semimartingale given by $X^{x}_{t}:=x+\int_{0}^{t}b_{s}^{o}ds+\int_{0}^{t}\sigma_{s}^{o}dB_{s}$ on the reference measure $\mathbb{P}$, where $(B_{s})_{s\in[0,T]}$ is a $d$-dimensional Brownian motion on $\mathbb{P}$, $(b_{s}^{o},\sigma_{s}^{o})_{s\in[0,T]}$ are baseline parameters, and $x\in\mathbb{R}^{d}$ is the initial state.
We then combine the penalization method of [21, 39, 54] (used to establish the well-posedness of reflected backward stochastic differential equations (BSDEs) characterizing (1.1)) with the entropy regularization framework of [66, 67] to propose and analyze the following optimal exploratory control problem under ambiguity:
(1.2) $\begin{aligned} \overline{V}_{t}^{x;N,\lambda}:=\operatorname*{ess\,sup}_{\pi\in\Pi}{\mathcal{}E}^{g}_{t}&[\int_{t}^{T}e^{-\int_{t}^{s}(\beta_{u}+N\pi_{u})du}\big(r(X_{s}^{x})+R(X_{s}^{x})\,N\pi_{s}-\lambda{\mathcal{}H}(\pi_{s})\big)\\ &\quad+e^{-\int_{t}^{T}(\beta_{u}+N\pi_{u})du}R(X_{T}^{x})],\end{aligned}$
where $\Pi$ is the set of all progressively measurable processes with values in $[0,1]$, representing Bernoulli-distributed controls randomizing stopping rules (see Remark 3.2), ${\mathcal{}H}:[0,1]\to\mathbb{R}$ denotes the binary differential entropy (see (3.1)), $\lambda>0$ represents the level of exploration to learn the unknown environment, and $N\in\mathbb{N}$ represents the penalization level (used for approximation of (1.1)).
In Theorem 3.4, we show that if $(b^{o},\sigma^{o})$ are sufficiently integrable (see Assumption 2), $r$ and $R$ has certain regularity and growth properties, and $\beta$ is uniformly bounded (see Assumption 2), then $\overline{V}^{x;N,\lambda}$ in (1.2) can be characterized by a solution of a BSDE. In particular, the optimal Bernoulli-distributed control of (1.2) is given by
(1.3) $\pi^{*,x;N,\lambda}_{t}:=\operatorname{logit}(\frac{N}{\lambda}(R(X_{t}^{x})-\overline{V}_{t}^{x;N,\lambda}))=[1+e^{-\frac{N}{\lambda}(R(X_{t}^{x})-\overline{V}_{t}^{x;N,\lambda})}]^{-1}$
@@
It is noteworthy that a similar logistic form as in (1.3) can also be observed in the non-robust setting in [15]; however, our value process $\overline{V}^{x;N,\lambda}$ is established through nonlinear expectation calculations. Moreover, the BSDE techniques of El Karoui et al. [21] are instrumental in the verification theorem for our maxmin problems (see Theorem 3.4). Lastly, our BSDE arguments enable a sensitivity analysis of $\overline{V}^{x;N,\lambda}$ with respect to the level of exploration; see Theorem 3.5 and Corollary 3.6.
@@
Next, under the same assumptions on $b^{o},\sigma^{o},r,R,\beta$, Theorem 4.1 establishes a policy iteration result. Specifically, at each step we evaluate the $g$-expectation value function under the control $\pi\in\Pi$ from the previous iteration and then update the control in the logistic form driven by this evaluated $g$-expectation value (as in (1.3)). This iterative process ensures that the resulting sequence of value functions and controls converge to the solution of (1.2) as the number of iterations goes to infinity.
@@
As an application of Theorem 4.1, under Markovian conditions on $b^{o},\sigma^{o},r,R,\beta$ (so that the assumptions made before hold), we devise an RL algorithm (see Algorithm 1) in which policy evaluation at each iteration, characterized by a PDE (see Corollary 4.3), can be implemented by the deep splitting method of Beck et al. [5].
@@
Finally, in order to illustrate all our theoretical results, we provide two numerical examples, American put-type and call-type stopping problems (see Section 5). We are able to observe policy improvement and convergence under several ambiguity degrees. Stability analysis for our exploratory BSDEs solution is also conducted with respect to ambiguity degree $\varepsilon$, temperature parameter $\lambda$ and penalty factor $N$ using put-type stopping problem, while robustness is shown by call-type stopping decision-making under different level of dividend rate misspecification.
@@ 1.1 Related literature
Sutton and Barto [63] opened up the field of RL, which has since gained significant attention, with successful applications [29, 44, 40, 60, 61]. In continuous-time settings, [66, 67] introduced an RL framework based on relaxed controls, motivating subsequent development of RL schemes [32, 35, 36, 37], applications and extensions [13, 14, 31, 64, 68].
Our formulation of exploratory stopping problems under ambiguity aligns with, and can be viewed as, a robust analog of [15, 17], who combine the penalization method for variational inequalities with the exploratory framework of [66, 67] in the PDE setting. Recently, an exploratory stopping-time framework based on a singular control formulation has also been proposed by [16].
While some proof techniques in our work bear similarities to those in [15, 17], the consideration of ambiguity introduces substantial differences. In particular, due to the Itô semimartingale setting of $X^{x}$ and the nonlinearity induced by the $g$-expectation, PDE-based arguments cannot be applied directly. Instead, we establish a robust (i.e., max–min) verification theorem using BSDE techniques. Building on this, we derive a policy iteration theorem by analyzing a priori estimates for iterative BSDEs. A related recent work of [26] proposes and analyzes an exploratory optimal stopping framework under discrete stopping times but without ambiguity. Lastly, we refer to [6, 7, 57] for machine learning (ML) approaches to optimal stopping.
Moving away from the continuous-time RL (or ML) results to the literature on continuous-time optimal stopping under ambiguity, we refer to [3, 4, 47, 51, 52, 58]. More recently, [43] proposes a framework for optimal stopping that incorporates both ambiguity and learning. Rather than adopting a worst-case approach, as in the above references, the framework employs the smooth ambiguity-aversion model of Klibanoff et al. [38] in combination with Bayesian learning.
@@ 1.2 Notations and preliminaries
  • $\mathbb{S}^{p}(\mathbb{R}^{k})$ is the set of all $\mathbb{R}^{k}$-valued, $\mathbb{F}$-progressively measurable càdlàg (i.e., right-continuous with left-limits) processes $Y=(Y_{t})_{t\in[0,T]}$ such that $\|Y\|_{\mathbb{S}^{p}}^{p}:=\mathbb{E}[\sup_{t\in[0,T]}|Y_{t}|^{p}]<\infty$;
@@ 2 Optimal stopping under ambiguity
Consider the optimal stopping time choice of an agent facing ambiguity, where the agent is ambiguity-averse and his/her stopping time is determined by observing an ambiguous underlying state process in a continuous-time environment. We model the agent’s preference and the environment by using the $g$-expectation ${\mathcal{}E}^{g}[\cdot]$ (see [12, 53]) defined as follows.
@@
    for every $(\omega,t)\in\Omega\times[0,T]$, $g(\omega,t,\cdot):\mathbb{R}^{d}\to\mathbb{R}$ is concave and $g(\omega,t,0)=0$.
Then we define ${\mathcal{}E}^{g}:L^{2}({\mathcal{}F}_{T};\mathbb{R})\ni\xi\to{\mathcal{}E}^{g}[\xi]\in\mathbb{R}$ as ${\mathcal{}E}^{g}[\xi]:=Y_{0},$ where $(Y,Z)\in\mathbb{S}^{2}(\mathbb{R})\times\mathbb{L}^{2}(\mathbb{R}^{d})$ is the unique solution of the following BSDE (see [49, Theorem 3.1]):
$Y_{t}=\xi+\int_{t}^{T}g(s,Z_{s})ds-\int_{t}^{T}Z_{s}dB_{s},$
@@
For (sufficiently integrable) $\mathbb{F}$-predictable processes $(b_{s}^{o})_{s\in[0,T]}$ and $(\sigma_{s}^{o})_{s\in[0,T]}$ taking values in $\mathbb{R}^{d}$ and $\mathbb{R}^{d\times d}$ respectively, we consider an Itô $(\mathbb{F},\mathbb{P})$-semimartingale $X^{x}:=(X^{x}_{t})_{t\in[0,T]}$ given by
@@
Either one of the following conditions is sufficient for Assumption 2 to hold true [2, Lemma 2.3]:
@@
    In fact, both sufficient conditions given in Remark 2.3 ensure that Assumption 2 holds for all $p\geq 2$ (see [41, Theorems 2.3.1 and 2.4.1])
@@
(2.2) $\begin{aligned} &V_{t}^{x}:=\operatorname*{ess\,sup}_{\tau\in{\mathcal{}T}_{t}}{\mathcal{}E}^{g}_{t}[\operatorname{I}_{t}^{x;\tau}];\qquad\operatorname{I}_{t}^{x;\tau}:=\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}r(X_{s}^{x})ds+e^{-\int_{t}^{\tau}\beta_{u}du}R(X_{\tau}^{x}),\end{aligned}$
@@
Let us that the $V^{x}$ given in (2.2) corresponds to a reflected BSDE with a lower obstacle. To that end, set for every $(\omega,t,y,z)\in\Omega\times[0,T]\times\mathbb{R}\times\mathbb{R}^{d}$ by
@@
Denote by $(Y_{t}^{x},Z_{t}^{x},K^{x}_{t})_{t\in[0,T]}$ a triplet of processes satisfying that
(2.4) $Y_{t}^{x}=R(X_{T}^{x})+\int_{t}^{T}F_{s}^{x}(Y_{s}^{x},Z_{s}^{x})ds-\int_{t}^{T}Z_{s}^{x}dB_{s}+K_{T}^{x}-K_{t}^{x},\;\;\mbox{for}\;t\in[0,T],$
We then introduce the notion of the reflected BSDE (see [39, Definition 2.1]). For this, recall the sets $\mathbb{S}^{2}(\mathbb{R})$ and $\mathbb{L}^{2}(\mathbb{R}^{d})$ given in Section 1.2.
@@
A triplet $(Y^{x},Z^{x},K^{x})$ is said to be a solution to the reflected BSDE (2.4) with the lower obstacle $(R(X_{t}^{x}))_{t\in[0,T]}$ if the following conditions hold:
@@
    $Y^{x}\in\mathbb{S}^{2}(\mathbb{R})$, $Z^{x}\in\mathbb{L}^{2}(\mathbb{R}^{d})$ and $K^{x}\in\mathbb{S}^{2}(\mathbb{R})$ which is nondecreasing and starts with $K_{0}^{x}=0$. Moreover, $(Y^{x},Z^{x},K^{x})$ satisfies (2.4);
@@
Under Assumptions 2 and 2, there exists a unique solution $(Y_{t}^{x}$, $Z_{t}^{x},K^{x}_{t})_{t\in[0,T]}$ of the reflected BSDE (2.4) with the lower obstacle $(R(X_{t}^{x}))_{t\in[0,T]}$ (see Definition 2.6). Indeed, one can easily show that the parameters of the reflected BSDE satisfy the conditions (i)–(iii) given in [39, Section 2], which enables to apply [39, Theorem 3.3] to ensures its existence and uniqueness to hold.
The following proposition establishes that the solution to the reflected BSDE (2.4) coincides with the Snell envelope of the optimal stopping problem under ambiguity given in (2.2). This result can be seen as a robust analogue of [20, Proposition 2.3] and [39, Proposition 3.1]. Several properties of (conditional) $g$-expectation developed in [12] are useful in the proof presented in Section 6.1.

### v2

@@ Abstract
We propose and analyze a continuous-time robust reinforcement learning framework for optimal stopping under ambiguity. In this framework, an agent chooses a robust exploratory stopping time motivated by two objectives: robust decision-making under ambiguity and learning about the unknown environment. Here, ambiguity refers to considering multiple probability measures dominated by a reference measure, reflecting the agent’s awareness that the reference measure representing her learned belief about the environment would be erroneous. Using the $g$-expectation framework, we reformulate the optimal stopping problem under ambiguity as a robust exploratory control problem with Bernoulli distributed controls. We then characterize the optimal Bernoulli distributed control via backward stochastic differential equations and, based on this, construct the robust exploratory stopping time that approximates the optimal stopping time under ambiguity. Last, we establish a policy iteration theorem and implement it as a reinforcement learning algorithm. Numerical experiments demonstrate the convergence, robustness, and scalability of our reinforcement learning algorithm across different levels of ambiguity and exploration.
@@
optimal stopping, exploratory control, reinforcement learning, ambiguity, robust optimization, $g$-expectation, policy iteration.
@@ 1 Introduction
Optimal stopping is a class of decision problems in which one seeks to choose a time to take a certain action so as to maximize an expected reward. It is applied in various fields, for instance to analyze the optimality of the sequential probability ratio test in statistics (e.g., [wald1948optimum]), to study consumption habits in economics (e.g., [dybvig1995dusenberry]) and notably to derive American option pricing (e.g., [peskir2006optimal]). A common challenge arising in all these fields is finding the best model to describe the underlying process or probability measure, which is usually unknown. Although significant efforts have been made to propose and analyze general stochastic models with improved estimation techniques, a margin of error in estimation inherently exists.
@@
In response to such model misspecification and estimation errors, recent works, Dai et al. [dai2024learning] and Dong [DongRLstopping], have cast optimal stopping problems within the continuous time reinforcement learning (RL) framework of Wang et al. [wang2020reinforcement] and Wang and Zhou [wang2020continuous]. Arguably, the exploratory (or randomized) optimal stopping framework is viewed as model-free, since agents, even without knowledge of the true model or underlying dynamics of the environment, can learn from observed data and determine a stopping rule that yields the best outcome. In this sense, the framework provides a systematic way to balance exploration and exploitation in optimal stopping.
Although model-free RL is sometimes interpreted as enabling direct deployment without an explicit model, this view can be misleading in practice. Limited and noisy training data may lead to inaccurate learned dynamics or value surrogates, and environmental conditions can change between training and execution; in physical or safety-critical systems, such discrepancies can entail severe consequences. Moreover, model-free learning is typically data-intensive, so one often relies on simulators or limited real data, thereby reintroducing ambiguity (or Knightian uncertainty) and creating a trade-off between safety during training and generalization at deployment [christiano2016transfer, morimoto2005robust]. In addition, real-world training is costly, time-consuming, and potentially unsafe, which often restricts data collection to a narrow set of scenarios and may result in overfitting and poor generalization under distributional shifts [pinto2017robust]. These considerations motivate robust RL approaches that seek policies whose performance remains reliable under model ambiguity and related disturbances.
This motivation is further reinforced by the observation that learning need not eliminate ambiguity at the level of beliefs. In particular, as argued in [epstein2003recursive] and emphasized by [marinacci1999limit], the link between empirical frequencies and asymptotic (learned) beliefs can be weakened by the degree of ambiguity in an agent’s prior beliefs, so that ambiguity may persist even with extensive learning. Consequently, the reliability of purely model-free exploratory frameworks can be fundamentally limited under ambiguity and related forms of model misspecification. In this sense, robust RL is both practically desirable and conceptually natural. Arguably, worst-case or minimax formulations provide a canonical framework for robust RL [garcia2015comprehensive-safeRL].
@@
The aim of this article is to propose and analyze a continuous-time RL framework for optimal stopping under ambiguity. Our framework starts with revisiting an optimal stopping problem under $g$-expectation (see Coquet et al. [coquet2002filtration], Peng [peng1997backward]): Let ${\mathcal{}T}_{t}$ be the set of all stopping times with values in $[t,T]$. Denote by ${\mathcal{}E}_{t}^{g}[\cdot]$ the conditional $g$-expectation with a driver $g:\Omega\times[0,T]\times\mathbb{R}^{d}\to\mathbb{R}$ , representing a worst-case expectation (see Definition 2.1, Remark 2.2). Then, the optimal stopping problem is
@@
where $(\beta_{t})_{t\in[0,T]}$ is a discount rate, $r:\mathbb{R}^{d}\to\mathbb{R}$ and $R:\mathbb{R}^{d}\to\mathbb{R}$ are reward functions, and $(X_{t}^{x})_{t\in[0,T]}$ is an Itô-semimartingale with an initial state $x\in\mathbb{R}^{d}$; see (2.1).
We then combine the penalization method of [el1997backward, lepeltier2005penalization, peng2005smallest] (used to establish the well-posedness of reflected backward stochastic differential equations (BSDEs) characterizing (1.1)) with the entropy regularization framework of [wang2020reinforcement, wang2020continuous] to study the following robust exploratory control problem:
(1.2) $\begin{aligned} \overline{V}_{t}^{x;N,\lambda}:=\operatorname*{ess\,sup}_{\pi\in\Pi}{\mathcal{}E}^{g}_{t}&\bigg[\int_{t}^{T}e^{-\int_{t}^{s}(\beta_{u}+N\pi_{u})du}\big(r(X_{s}^{x})+R(X_{s}^{x})\,N\pi_{s}-\lambda{\mathcal{}H}(\pi_{s})\big)ds\\ &\quad+e^{-\int_{t}^{T}(\beta_{u}+N\pi_{u})du}R(X_{T}^{x})\bigg],\end{aligned}$
where $\Pi$ is the set of all $[0,1]$-valued, Bernoulli-distributed controls randomizing stopping rules; see Remark 3.2, ${\mathcal{}H}(a)$ denotes the binary differential entropy for $a\in[0,1]$; see (3.1), $\lambda>0$ represents the level of exploration, and $N\in\mathbb{N}$ represents the penalization level used for approximation of (1.1).
In Theorem 3.4, we show that if the characteristics of $(X_{t}^{x})_{t\in[0,T]}$ are sufficiently integrable (see Assumption 2), $r$ and $R$ satisfy certain regularity and growth conditions, and $(\beta_{t})_{t\in[0,T]}$ is uniformly bounded (see Assumption 2), then the robust exploratory control problem $\overline{V}^{x;N,\lambda}$ in (1.2) can be characterized by a BSDE (see (3.8)). Moreover, its optimal Bernoulli-distributed control is given by
(1.3) $\pi^{*,x;N,\lambda}_{t}:=\operatorname{logit}\bigg(\frac{N}{\lambda}(R(X_{t}^{x})-\overline{V}_{t}^{x;N,\lambda})\bigg)\in\Pi,$
@@
Next, using $\pi^{*,x;N,\lambda}$ in (1.3), we define a stopping time by
(1.4) ${\tau}^{x;N,\lambda}_{t}:=\inf\bigg\{s\geq t:\pi^{*,x;N,\lambda}_{s}\geq\frac{1}{1+2^{N(T-s)}}\bigg\}\wedge T\in{\mathcal{}T}_{t}.$
Theorem 3.6 shows that for any sequences $(N_{n},\lambda_{n})_{n\in\mathbb{N}}$ such that $N_{n}\uparrow\infty$ and $\lambda_{n}\downarrow 0$ as $n\to\infty$, ${\tau}^{x;N_{n},\lambda_{n}}_{t}$ converges to the optimal stopping time of $V^{x}_{t}$ in (1.1). Moreover, the corresponding $g$-expectation value converges to $V_{t}^{x}$. In other words, under sufficiently large $N$ and sufficiently small $\lambda$, the stopping time ${\tau}^{x;N,\lambda}_{t}$ in (1.4), induced by the optimal Bernoulli-distributed control in (1.3), approximates the optimal stopping time of $V_{t}^{x}$. Hence, we refer to ${\tau}^{x;N,\lambda}_{t}$ as a robust exploratory stopping time.
It is noteworthy that a logistic form for exploratory controls as in (1.3) already appears in the non-robust exploratory stopping problem in [dai2024learning]. However, the natural question—namely, to what extent the optimal stopping time of the original stopping problem can be approximated from the exploratory controls—remains open therein. Theorem 3.6 provides an answer to this question, even in a robust framework. In particular, our result also applies to the non-robust setting.
@@
Last, under the same assumptions on $X^{x},r,R$, and $\beta$, Theorem 4.1 establishes a policy iteration result. Specifically, at each step we evaluate the $g$-expectation value function under the control $\pi\in\Pi$ from the previous iteration and then update the control in the logistic form driven by this evaluated $g$-expectation value as in (1.3). This iterative process ensures that the resulting sequence of value functions and controls converge to the solution of (1.2) as the number of iterations goes to infinity.
@@
As an application of Theorem 4.1, we devise an RL algorithm (see Algorithm 1) in which policy evaluation at each iteration, characterized by a PDE (see Corollary 4.5), can be implemented by the deep splitting method of Beck et al. [Becker-Jentzen-Neufeld-Deep-Splitting-21].
@@
Using our RL algorithm, Section 5 presents several numerical examples, including robust American put- and call-option valuation. In the one-dimensional setting, we observe policy improvement and convergence under several ambiguity degrees. Moreover, stability analysis for our exploratory BSDEs solution is conducted with respect to ambiguity degree $\varepsilon$, exploration level $\lambda$ and penalization level $N$ in the put-type stopping problem, while robustness is illustrated in the call-type problem under different levels of dividend rate misspecification. Finally, we further investigate the policy improvement and convergence for a put-type Basket option valuation, and robustness for a call-type geometric option in multi-dimensional settings, thereby supporting the scalability of our algorithm.
Beyond the valuation of American options, our robust exploratory stopping framework and RL algorithm can be applied in a broad class of optimal stopping problems under ambiguity, for instance to determine early termination rules in clinical trials under unknown treatment response distributions (e.g., [pocock1977group]), to analyze Gittins-type indices in multi-armed bandit problems under unknown transition dynamics (e.g., [gittins2011multi]), and to inform maintenance decisions such as repair or replacement of deteriorating systems under imperfectly observed failure signals (e.g., [pierskalla1976survey]), among others.
Our main contributions can be summarized as follows. Beyond the standard BSDE formulation of the robust exploratory control framework, we uncover a structural link between robust exploratory controls and optimal stopping times under ambiguity, and prove—via a priori BSDE estimates and a convergence of monotone hitting times—that the robust exploratory stopping time approximates the optimal stopping time under ambiguity. Moreover, we establish a policy improvement theorem in a general setting that allows unbounded payoffs and arbitrary initial policy, while accommodating an ambiguity-induced driver in the generator; see Remarks 4.2, 4.6 for comparisons with related works [dai2024learning, DongRLstopping]. Finally, in our RL algorithm, we propose a new policy evaluation scheme based on the deep splitting method, combined with a quadratic variation–based procedure using the observed data to handle the driver term induced by the $g$-expectation; see (4.5) in Section 4. Numerical experiments demonstrate the convergence, robustness, and scalability of the proposed algorithm.
@@ 1.1 Related literature
Sutton and Barto [SuttonBarto1998] opened up the field of RL which has since gained significant attention with notable applications [Guo2025DeepSeekR1-nature, Mnih2015HumanLevel, autodriver, Silver2016Mastering, Silver2017MasteringWithout]. In continuous-time settings, [wang2020reinforcement, wang2020continuous] introduced an RL framework based on exploratory or relaxed controls, motivating subsequent development of RL schemes [huang2025convergence, Policy-evaluation, Policy-gradient, small-q], applications and extensions [DaiLearning2023, DaiRecursive20203, han2023choquet, exploratoryHJB, Liwu2024]. A substantial literature on robust RL exists in a discrete-time setting [blanchet2023double, panaganti2022robust, RobustRL-model-mismatch, RobustRL-deep], while a continuous-time robust RL remains less explored.
Our robust exploratory stopping problem under ambiguity aligns with, and can be viewed as, a robust analog of [dai2024learning, DongRLstopping]. Recently, an exploratory stopping-time framework based on a singular control formulation has also been proposed by [dianetti2024exploratory]. A related recent work of [frikha2025entropy] studies an exploratory optimal stopping framework under discrete stopping times but without ambiguity. Lastly, we refer to [becker2019deep, becker2021solving, reppen2025neural] for machine learning (ML) approaches to optimal stopping.
Moving away from the continuous-time RL (or ML) results to the literature on continuous-time optimal stopping under ambiguity, we refer to [bayraktar2011optimal_a, bayraktar2011optimal_b, nutz2015optimal, park2023irreversible, PW23, riedel2009optimal]. More recently, [tankov2025optimalstoppingdivestmenttiming] proposes a framework for optimal stopping that incorporates both ambiguity and learning. Rather than adopting a worst-case approach, it employs the smooth ambiguity framework of Klibanoff et al. [klibanoff2005smooth] and the Bayesian learning.
@@ 1.2 Notations and preliminaries
  • $\mathbb{S}^{p}(\mathbb{R}^{k})$ is the set of all $\mathbb{R}^{k}$-valued, $\mathbb{F}$-progressively measurable processes $Y=(Y_{t})_{t\in[0,T]}$ such that $\|Y\|_{\mathbb{S}^{p}}^{p}:=\mathbb{E}[\sup_{t\in[0,T]}|Y_{t}|^{p}]<\infty$;
@@ 2 Optimal stopping under ambiguity
Consider the optimal stopping of an agent facing ambiguity, where the agent is ambiguity-averse and his/her stopping time is determined by observing an ambiguous underlying state process in a continuous-time environment. We model the agent’s preference and the environment by using the $g$-expectation ${\mathcal{}E}^{g}[\cdot]$ (see [coquet2002filtration, peng1997backward]) defined as follows.
@@
    for every $(\omega,t)\in\Omega\times[0,T]$, $g(\omega,t,\cdot):\mathbb{R}^{d}\to\mathbb{R}$ is concave, $g(\omega,t,0)=0$, and satisfies that $g(\omega,t,cz)=cg(\omega,t,z)$ for all $z\in\mathbb{R}^{d}$ and $c\in(0,1)$.
Then we define ${\mathcal{}E}^{g}:L^{2}({\mathcal{}F}_{T};\mathbb{R})\ni\xi\to{\mathcal{}E}^{g}[\xi]\in\mathbb{R}$ as ${\mathcal{}E}^{g}[\xi]:=Y_{0},$ where $(Y,Z)\in\mathbb{S}^{2}(\mathbb{R})\times\mathbb{L}^{2}(\mathbb{R}^{d})$ is the unique solution of the following BSDE (see [pardoux1990adapted, Theorem 3.1]):
$Y_{t}=\xi+\int_{t}^{T}g(s,Z_{s})ds-\int_{t}^{T}Z_{s}dB_{s},\quad t\in[0,T],$
@@
For (sufficiently integrable) $\mathbb{F}$-predictable processes $(b_{s}^{o})_{s\in[0,T]}$ and $(\sigma_{s}^{o})_{s\in[0,T]}$ taking values in $\mathbb{R}^{d}$ and $\mathbb{R}^{d\times d}$ respectively, we consider an Itô $(\mathbb{F},\mathbb{P})$-semimartingale
@@
Either one of the following conditions is sufficient for Assumption 2 to hold true [bartl2023sensitivity, Lemma 2.3]:
@@
    In fact, both sufficient conditions given in Remark 2.3 ensure that Assumption 2 holds for all $p\geq 2$ (see [mao2007stochastic, Theorems 2.3.1 and 2.4.1])
@@
(2.2) $\begin{aligned} &V_{t}^{x}:=\operatorname*{ess\,sup}_{\tau\in{\mathcal{}T}_{t}}{\mathcal{}E}^{g}_{t}[\operatorname{I}_{t}^{x;\tau}],\quad\operatorname{I}_{t}^{x;\tau}:=\int_{t}^{\tau}e^{-\int_{t}^{s}\beta_{u}du}r(X_{s}^{x})ds+e^{-\int_{t}^{\tau}\beta_{u}du}R(X_{\tau}^{x}),\end{aligned}$
@@
In the following, we remark that $V^{x}$ given in (2.2) corresponds to a reflected BSDE with an obstacle. To this end, set for every $(\omega,t,y,z)\in\Omega\times[0,T]\times\mathbb{R}\times\mathbb{R}^{d}$ by
@@
Denote by $(Y^{x},Z^{x},K^{x})$ a triplet of processes satisfying that
(2.4) $Y_{t}^{x}=R(X_{T}^{x})+\int_{t}^{T}F_{s}^{x}(Y_{s}^{x},Z_{s}^{x})ds-\int_{t}^{T}Z_{s}^{x}dB_{s}+K_{T}^{x}-K_{t}^{x},\quad t\in[0,T].$
We then introduce the notion of the reflected BSDE (see [el1997reflected, Section 2]). For this, recall the sets $\mathbb{S}^{2}(\mathbb{R})$ and $\mathbb{L}^{2}(\mathbb{R}^{d})$ given in Section 1.2.
@@
A triplet $(Y^{x},Z^{x},K^{x})$ is said to be a solution to the reflected BSDE (2.4) with the obstacle $(R(X_{t}^{x}))_{t\in[0,T]}$ if the following conditions hold:
@@
    $Y^{x}\in\mathbb{S}^{2}(\mathbb{R})$, $Z^{x}\in\mathbb{L}^{2}(\mathbb{R}^{d})$, and $K^{x}\in\mathbb{S}^{2}(\mathbb{R})$ is a nondecreasing, continuous process with $K_{0}^{x}=0$. Moreover, $(Y^{x},Z^{x},K^{x})$ satisfies (2.4);
@@
Under Assumptions 2 and 2, the parameters of the reflected BSDE (2.4) with the obstacle $(R(X_{t}^{x}))_{t\in[0,T]}$ satisfy the conditions (i)–(iv) given in [el1997reflected, Section 2]. This enables to apply [el1997reflected, Theorem 5.2] to ensure that there exists a unique solution $(Y^{x},Z^{x},K^{x})$ of the reflected BSDE; see Definition 2.6.
The following proposition establishes that the solution to the reflected BSDE (2.4) coincides with the Snell envelope of the optimal stopping problem under ambiguity given in (2.2). This result can be seen as a robust analogue of [el1997reflected, Proposition 2.3] and [lepeltier2005penalization, Proposition 3.1]. Several properties of (conditional) $g$-expectation developed in [coquet2002filtration] are useful in the proof presented in Section 6.1.